At s.t.p,a mass of gas occupies a volume of of 802cm³.find the pressure at which the gas occupies 950cm³at 320k


1
Expert's answer
2022-02-09T01:42:09-0500

pV = nRT

STP: T = 273 K, p = 1 atm

R = 8.205 m3 atm K−1 mol−1

n = pV / RT = (1 x 0.802) / (8.205 x 273) = 0.0004 mol

p2 = nRT / V = (0.0004 x 8.205 x 302) / 0.905 = 0.98 atm